A container with a closed pinhole contains 1 mol of H2(g) and 1 mol of He(g). When the pinhole is opened, H2(g) starts to effuse at a rate of 0.48 mol/h. What is the rate of effusion of He(g) at that moment?

  •  A. 1.21 mol/h
  •  B. 0.34 mol/h
  •  C. 0.68 mol/h
  •  D. 0.24 mol/h
  •  E. 0.95 mol/h
1
Expert's answer
2021-03-08T05:52:41-0500

Solution:

Graham's law of effusion:



From Graham’s law, we have:

(rate of effusion of H2) / (rate of effusion of He) = (MHe)1/2 / (MH2)1/2

The molar mass of He is 4.0026 g/mol.

The molar mass of H2 is 2.016 g/mol.

Hence,

(0.48) / (rate of effusion of He) = (4.0026)1/2 / (2.016)1/2

(0.48) / (rate of effusion of He) = (2.00065) / (1.41986)

rate of effusion of He = 0.48 × 1.41986 / 2.00065 = 0.34065 = 0.34

rate of effusion of He = 0.34 mol/h


Answer: B. 0.34 mol/h.